OpenStudy (anonymous):

examples of funtions for: 1. lim x to infinity f(x)/g(x) = infinity 2. lim x to infinity f(x)/g(x) = 6 3. lim x to infinity f(x)/g(x) = 0

OpenStudy (anonymous):

\[\lim_{x\to \infty}\frac{f(x)}{g(x)}=\infty\] take any polynomial for the numerator whose degree is larger than the degree of the denominator

OpenStudy (anonymous):

you can use \[\frac{x^2+2}{x}\] or anything of your choosing

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty} \frac{f(x)}{g(x)}=6; \lim_{x \rightarrow \infty} \frac{6x+7}{x+9}=6\]